OpenStudy (anonymous):

use the multiplacation principle to solve each equation. -x = 7/13

OpenStudy (anonymous):

Multiply both sides of the equation by -1, you get -1(-x) = -1(7/13) Therefore, x = -7/13. The multiplication Principle basically says that you can multiple both sides of an equation by the same number. In this problem, we multiplied both sides by -1 so as to have x alone, not -x.